OpenStudy (anonymous):

check whether the given number is a solution of the equation or inequality. 2p-1 >7;3 -

OpenStudy (anonymous):

\[2p-1\ge 7;3\]

OpenStudy (helder_edwin):

this is the same as your previous post. just replace p=3 then u have \[ \large 2p-1=2\cdot3-1=6-1=5 \] is 5>=7 ??

OpenStudy (anonymous):

\[P=5\ge 7\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

soit is not an inequality?

OpenStudy (anonymous):

??????????????????????